# An exact approach for the multi-depot electric vehicle scheduling problem

### v1

@@ 4.1 A 3-index formulation
We introduce two sets of decision variables for this 3-index formulation: For each $k\in K$ and each $(i,j)\in A$ let $x_{ij}^{k}\in\{0,1\}$ be a binary variable, which is 1, if the arc $(i,j)$ is used from vehicles associated with depot $k$, 0 otherwise. Furthermore, for each $i\in V$, let $\varepsilon_{i}$ be the energy level of vehicles when leaving node $i$. With these decision variables the 3-index MDEVSP can be formulated as the mixed-integer multi-commodity flow model
$\min\leavevmode\nobreak\ w_{1}\Big{(}\sum_{k\in K}\sum_{j\in A^{+}(o_{k})}$ $x^{k}_{o_{k},j}\Big{)}+w_{2}\Big{(}\sum_{k\in K}\sum_{c\in V^{C}}\sum_{i\in A^{-}(c)}x^{k}_{ic}\Big{)}+$ $w_{3}\Big{(}\sum_{k\in K}\sum_{(i,j)\in A}p_{ij}x^{k}_{ij}\Big{)}$ (1a)
[…]
@@
Constraints (1h) set the energy level at all origin depots to the maximum energy level (all vehicles leave the depots fully charged). Constraints (1i) make sure that the energy level when leaving any node except the depots is not below the minimum required energy level. Constraints (1j) ensure that the energy level when returning to the depot is not below the required energy level at the destination depot. Constraints (1k) make sure that the energy level is not below the minimum required energy level before arriving at the charging station (which is implied by the fact that the energy level is not below the minimum required energy level when arriving at the next node after a service trip node, as charging nodes are always preceded by service trip nodes). Constraints (1l) guarantee that the energy level at the end of a service trip corresponds to the energy level of the previous node minus the energy that was consumed by the deadhead trip connecting the previous node and this service trip and the energy used by the service trip itself. Constraints (1m) allow the energy level at the end of a charging node to rise to at most the energy level of the previous node minus the energy that was consumed by the deadhead trip connecting the previous node and the charging station plus the energy that can be loaded at this charging node. Constraints (1n) make sure that the energy level at the destination depots correspond to at most the energy level at the previous node minus the energy spent on the deadhead trip connecting the previous node to the depot. Constraints (1o) make sure that a vehicle cannot be charged to an energy level that is higher than the battery capacity. The domain of the variable $x$ is defined in (1p).
Note, that (1) is a compact mixed-integer linear program, that can be solved with off-the-shelve solvers. We will present computational results for doing so later on.
@@ 4.2 A 2-index formulation
$\min\leavevmode\nobreak\ w_{1}\Big{(}\sum_{k\in K}\sum_{j\in A^{+}(o_{k})}$ $x_{o_{k},j}\Big{)}+w_{2}\Big{(}\sum_{c\in V^{C}}\sum_{i\in A^{-}(c)}x_{ic}\Big{)}+$ $w_{3}\Big{(}\sum_{(i,j)\in A}p_{ij}x_{ij}\Big{)}$ (2a)
[…]
@@
All other constraints of the 2-index formulation (2) are defined analogously to the constraints of the 3-index formulation (1), with the difference that now, we do not have a separate variable $x^{k}_{ij}$ for each $k\in K$ representing vehicles associated to the depot $k\in K$, but only one variable $x_{ij}$ for each arc $(i,j)\in A$. Thus, it is not possible to tell directly from the variables of (2) at which depot a vehicle using the arc $(i,j)$ departed. This advantage of having fewer variables comes at the cost of having a high number of infeasible path constraints (2f). We describe in Section 5.1 how to deal with this computationally.
Another option to avoid the undesirable infeasible paths is by replacing constraints (2f) with other constraints similar to how it is done, e.g. in Parragh, (2011). For doing so, for each $k\in K$ we introduce the set $\mathcal{U}_{k}$ of all node subsets $U\subseteq V$, such that the origin depot $o_{k}\in U$ and the destination depot $d_{k}\notin U$, while for all other $k^{\prime}\in K\setminus\{k\}$ the origin depots $o_{k^{\prime}}\notin U$ and the destination depots $d_{k^{\prime}}\in U$. Then, we can replace the infeasible path constraints (2f) in the 2-index formulation (2) with
$\sum_{i\in U}\leavevmode\nobreak\ \sum_{j\in A^{+}(i),j\notin U}x_{ij}\geq\sum_{j\in A^{+}(o_{k})}x_{o_{k},j}\qquad\forall k\in K,\,U\in\mathcal{U}_{k}.$ (3)
@@ 4.4 Extension to alternative zero emission technologies
For example, our model is able to depict the possibility of opportunity charging at the start or end locations of the service trips. To do so, for each service trip $i\in V^{I}$, if opportunity charging is available at the start (end) location of the service trip $\ell_{i}^{s}$ ($\ell_{i}^{e}$), then a charging station $a$ is added to $C$ at location $\ell_{a}=\ell_{i}^{s}$ ($\ell_{a}=\ell_{i}^{e}$) and the corresponding values of $t_{ij}$, $d_{ij}$ and $p_{ij}$ for $i=a$ or $j=a$ need to be adapted accordingly. Only then the graph is constructed.
@@
Furthermore, fuel cell electric buses can be considered with our methodology if the charging stations $a\in C$ represent hydrogen fueling stations, and the energy consumption rate $\theta$, the charging rate $r$, the maximum energy level $s^{max}$, the minimum energy level $s^{min}$ and $s^{min}_{dep}$ and the minimum time that needs to be available for charging $t^{min}$ are modified to fit for the fuel cell case.
Thus, our approach is universal in the sense that it can be adapted to various zero emission technology settings. Also, the consideration of diesel buses is possible with our model analogously to the fuel cell case.
@@ 5 Branch-and-cut algorithm
In order to employ the previously introduced 3-index formulation (1) and 2-index formulation (2) for solving the MDEVSP, one could use one of many available MILP solvers in the standard configuration. While this is possible for (1), the large number of infeasible path constraints (2f) or connectivity constraints (3) becomes prohibitive for (2).
Thus, we have developed a branch-and-cut algorithm for (2). Branch-and-cut algorithms incorporate the principles of branch-and-bound and pair it with the cutting-plane idea. They start from solving the linear relaxation of the MILP, while considering only a reasonable subset of the original constraints. Typically, constraints of exponential size are excluded. Then, in the course of the algorithm a separation method is required, which finds violated constraints and adds them in an iterative fashion, until ultimately no original constraints are violated anymore, even though they might not be included explicitly.
@@ 6 Computational experiments
We are now able to present computational results. In the following, first, the attributes of the generated test instances are described in Section 6.1. Then, our obtained results are discussed for instances with a single depot in Section 6.2, a low number of depots (two, three, and four) in Section 6.3 and a high number of depots (six and eight) in Section 6.4.
@@ 6.1 Computational setup and test instances
Everything is implemented in Julia 1.11.1. For the branch-and-cut algorithms CPLEX 22.1 is employed, where for the separation as described in Section 5 we utilize LazyConstraint for separating integer solutions and UserCut for separating fractional solutions. In the separation of the connectivity constraints (3) described in Section 5.2 we use the Boykov-Kolmogorov algorithm within the function maximum_flow() of Julia in order to determine a minimum cut. All experiments were carried out on a Quad-core X5570 Xeon CPU @2.93GHz with a memory of 48 GB.
@@
For our computational experiments, we use a set of MDEVSP benchmark instances generated in a similar way to the generation of MDVSP class A instances in Carpaneto et al., (1989) used by many other researchers (see Gkiotsalitis et al., (2023); Bianco et al., (1994); Fischetti et al., (2001); Forbes et al., (1994); Ribeiro and Soumis, (1994)). The instances of Carpaneto et al., (1989) are not directly applicable to our setting, as they do not involve electric vehicles, which necessitate the inclusion of charging infrastructure in the instance data. Thus, we have created instances as close as possible to Carpaneto et al., (1989) and adapted them to the electric case as it was done in Gkiotsalitis et al., (2023).
@@
In particular, we consider instances with $|V^{I}|\in\{10,20,30,40,50,60\}$ service trips, $|K|=|O|=|D|\in\{1,2,3,4,6,8\}$ depots, and $|C|\in\{1,2,3\}$ charging stations. For each instance, we first determine a number $\nu$ of potential start and end locations of service trips (so-called relief locations) by choosing $\nu$ as uniformly random integer in $\left[\frac{1}{3}|V^{I}|,\frac{1}{2}|V^{I}|\right]$. Then we choose the $\nu$ relief locations $\ell^{ST}_{1}$, …, $\ell^{ST}_{\nu}$, the depot locations $\ell_{k}$ for each $k\in K$, as well as the charging station locations $\ell_{a}$ for each $a\in C$ randomly distributed in a 60 km by 60 km square in the Euclidean plane using a uniform distribution, resulting in coordinates (latitude and longitude) for each location. Finally, for each service trip $i\in V^{I}$ we decide if it is a short trip (with probability $40\%$, representing urban journeys) or a long trip (with probability $60\%$, representing extra-urban journeys that start and end at the same location). Then for each service trip $i\in V^{I}$ we choose both the start and the end location of the service trip $\ell^{s}_{i}$ and $\ell^{e}_{i}$ uniformly at random as one of the $\nu$ relief locations $\ell^{ST}_{1}$, …, $\ell^{ST}_{\nu}$, where we make sure that $\ell^{s}_{i}=\ell^{e}_{i}$ holds for long service trips $i$. From these locations we compute the distances $d_{ij}$ as the Euclidean distances between the (end) location of $i$ and the (start) location of $j$ for each $i,j\in V^{I}\cup O\cup D\cup C$. For the travel times $t_{ij}$ (in minutes) we assume an average vehicle speed of 60 km per hour translating to 1 km per minute, so $t_{ij}=d_{ij}$.
@@
For each short service trip $i\in V^{I}$ we generate the start time $s_{i}$ (in minutes since midnight) as a random integer in the interval [420,480] with a probability of 15%, in the interval [480,1020] with a probability of 70% and in the interval [1020,1080] with a probability of 15%. Moreover, we choose the end time $e_{i}$ uniformly at random as integer in the interval $[s_{i}+d_{i}+5,s_{i}+d_{i}+40]$, where $d_{i}$ is the Euclidean distance between $\ell^{s}_{i}$ and $\ell^{e}_{i}$. For each long service trip $i\in V^{I}$ we generate the start time $s_{i}$ as a uniform random integer in the interval [300,1200] and the end time $e_{i}$ as a uniform random integer in the interval $[s_{i}+180,s_{i}+300]$. For each service trip $i\in V^{I}$ this yields the duration $u_{i}$, from which we compute the energy usage $q_{i}$ as $q_{i}=\theta u_{i}$.
@@
For each depot index $k\in K$ we generate the number of vehicles available as $b_{k}$ as a uniformly random integer in $[3+\frac{1}{3|K|}|V^{I}|,3+\frac{1}{2|K|}|V^{I}|]$. The parameters related to vehicles and charging are set to $\theta=1.3$, $s^{max}=1000$, $s^{min}=10$ and $r=\frac{50}{6}$ as it is done in Gkiotsalitis et al., (2023). Note, that this yields $t^{max}=\frac{1}{r}(s^{max}-s^{min})=118.8$. Furthermore, we use $s^{min}_{dep}=0.7s^{max}$ and $t^{min}=\frac{1}{100}s^{max}$.
@@
For each instance, we first generate the graph $G=(V,A)$ as described in Section 3.2. As this can be done quite fast (2.35 seconds for the largest instance) we omit the running time for doing so from now on. This yields graphs with a broad range of sizes, ranging from $|V|=40$ and $|A|=190$ for one of the smallest instances with $|V^{I}|=10$, $|K|=1$ and $|C|=1$ up to $|V|=917$ and $|A|=9120$ for one of the largest instances with $|V^{I}|=60$, $|K|=8$ and $|C|=3$. Note that the majority of the nodes of $G$ is charging nodes ($|V^{C}|=851$ for the latter instance with $|V|=917$).
@@
The graphs $G=(V,A)$ enable us to utilize the 3-index formulation (1) and the 2-index formulation (2) and its variant. For the weights of the objective functions of (1) and (2) we carefully engineered $w_{1}=100,000$, $w_{2}=4,000$ and $w_{3}=1$ as appropriate weights for all our instances, to make sure that we indeed minimize the desired quantities (number of vehicles, number of charging events, energy usage for deadhead trips) in lexicographic order.
@@ 6.2 Results for a single depot
Table 2: Results for a single depot
$|V^{I}|$ $|C|$opt $t$(s) gap(%) #nB&B $z^{*}$ $LB_{R}$
10 1 10 0.05 0.00 22.6 471,689.05 469,066.54
10 2 10 0.07 0.00 3.2 471,093.91 469,758.28
10 3 10 0.30 0.00 0.0 552,536.49 551,906.75
20 1 10 0.11 0.00 178.4 864,496.71 862,890.77
20 2 10 0.24 0.00 217.4 843,637.16 841,021.47
20 3 10 0.68 0.00 864.0 760,955.41 756,186.32
30 1 10 1.54 0.00 1,910.1 1,112,737.81 1,100,675.44
30 2 10 0.95 0.00 879.0 1,154,218.11 1,146,477.21
30 3 10 3.02 0.00 1,624.5 1,155,001.82 1,141,261.63
40 1 10 6.53 0.00 5,111.9 1,466,963.22 1,450,335.05
40 2 10 321.87 0.00 209,772.8 1,486,804.98 1,467,894.73
40 3 8 2,167.82 0.30 846,177.5 1,464,963.46 1,444,291.56
50 1 10 70.63 0.00 35,872.8 1,613,935.37 1,572,182.94
50 2 9 2,497.99 0.21 942,411.9 1,734,256.53 1,693,949.86
50 3 8 2,174.41 0.94 743,048.8 1,654,094.08 1,624,816.48
60 1 8 4,388.87 0.19 1,327,319.6 2,043,139.67 1,998,699.10
60 2 8 2,164.61 0.20 365,917.0 2,208,161.83 2,167,048.28
60 3 8 4,062.17 0.44 944,193.8 1,962,828.15 1,932,322.04
In Table 2 we display results for a single depot. In particular, we present the number (#opt) of instances that were solved to optimality (out of 10), the average CPU times ($t$) in seconds, the average percentage gap at the time limit (3 hours) of all instances that were not solved to optimality, the average number of B&B nodes (#nB&B), the average best found objective function value $z^{*}$ at termination and the average lower bound at the root node $LB_{R}$.

### v2

@@ 4.1 A 3-index formulation
We introduce two sets of decision variables for this 3-index formulation: For each $k\in K$ and each $(i,j)\in A$ let $x_{ij}^{k}\in\{0,1\}$ be a binary variable, which is 1, if the arc $(i,j)$ is used from vehicles associated with depot $k$, 0 otherwise. Furthermore, for each $i\in V$, let $\varepsilon_{i}$ be the energy level of vehicles when leaving node $i$. With these decision variables, the 3-index MDEVSP can be formulated as the following mixed-integer multi-commodity flow model.
$\min\penalty 10000\ w_{1}\Big(\sum_{k\in K}\sum_{j\in A^{+}(o_{k})}$ $x^{k}_{o_{k},j}\Big)+w_{2}\Big(\sum_{k\in K}\sum_{c\in V^{C}}\sum_{i\in A^{-}(c)}x^{k}_{ic}\Big)+$ $w_{3}\Big(\sum_{k\in K}\sum_{(i,j)\in A}p_{ij}x^{k}_{ij}\Big)$ (1a)
$\text{s.t.\penalty 10000\ }\sum_{k\in K}\sum_{i\in A^{-}(j)}x_{ij}^{k}$ $=1$ $\forall j\in V^{I}$ (1b)
@@
Constraints (1h) set the energy level at all origin depots to the maximum energy level (all vehicles leave the depots fully charged). Constraints (1i) make sure that the energy level when leaving any node except the depots is not below the minimum required energy level. Constraints (1j) ensure that the energy level when returning to the depot is not below the required energy level at the destination depot. Constraints (1k) make sure that the energy level is not below the minimum required energy level before arriving at the charging station (which is implied by the fact that the energy level is not below the minimum required energy level when arriving at the next node after a service trip node, as charging nodes are always preceded by service trip nodes). Constraints (1l) guarantee that the energy level at the end of a service trip corresponds to the energy level of the previous node minus the energy consumed by the deadhead trip connecting the previous node and this service trip and the energy used by the service trip itself. Constraints (1m) allow the energy level at the end of a charging node to rise to at most the energy level of the previous node minus the energy that was consumed by the deadhead trip connecting the previous node and the charging station plus the energy that can be loaded at this charging node. Constraints (1n) make sure that the energy level at the destination depots corresponds to at most the energy level at the previous node minus the energy spent on the deadhead trip connecting the previous node to the depot. Constraints (1o) make sure that a vehicle cannot be charged to an energy level that is higher than the battery capacity. The domain of the variable $x$ is defined in (1p).
Note that (1) is a compact mixed-integer linear program, that can be solved with off-the-shelf solvers. We will present computational results for doing so later on.
@@ 4.2 A 2-index formulation
$\min\penalty 10000\ w_{1}\Big(\sum_{k\in K}\sum_{j\in A^{+}(o_{k})}$ $x_{o_{k},j}\Big)+w_{2}\Big(\sum_{c\in V^{C}}\sum_{i\in A^{-}(c)}x_{ic}\Big)+$ $w_{3}\Big(\sum_{(i,j)\in A}p_{ij}x_{ij}\Big)$ (2a)
$\text{s.t.\penalty 10000\ }\sum_{i\in A^{-}(j)}x_{ij}$ $=1$ $\forall j\in V^{I}$ (2b)
@@
All other constraints of the 2-index formulation (2) are defined analogously to the constraints of the 3-index formulation (1), with the difference that now we do not have a separate variable $x^{k}_{ij}$ for each $k\in K$ representing vehicles associated to the depot $k\in K$, but only one variable $x_{ij}$ for each arc $(i,j)\in A$. Thus, it is not possible to directly tell from the variables of (2) at which depot a vehicle using the arc $(i,j)$ departed. This advantage of having fewer variables comes at the cost of having a high number of infeasible path constraints (2f). We describe in Section 5.1 how to deal with this computationally.
Another option to avoid the undesirable infeasible paths is to replace constraints (2f) with other constraints similar to how it is done, e.g., in Parragh, (2011). For doing so, for each $k\in K$ we introduce the set $\mathcal{U}_{k}$ of all node subsets $U\subseteq V$, such that the origin depot $o_{k}\in U$ and the destination depot $d_{k}\notin U$, while for all other $k^{\prime}\in K\setminus\{k\}$ the origin depots $o_{k^{\prime}}\notin U$ and the destination depots $d_{k^{\prime}}\in U$. Then, we can replace the infeasible path constraints (2f) in the 2-index formulation (2) with
$\sum_{i\in U}\penalty 10000\ \sum_{j\in A^{+}(i),j\notin U}x_{ij}\geq\sum_{j\in A^{+}(o_{k})}x_{o_{k},j}\qquad\forall k\in K,\,U\in\mathcal{U}_{k}.$ (3)
@@ 4.4 Extension to alternative zero emission technologies
For example, our model is able to depict the possibility of opportunity charging at the start or end locations of service trips. To do so, for each service trip $i\in V^{I}$, if opportunity charging is available at the start (end) location of the service trip $\ell_{i}^{s}$ ($\ell_{i}^{e}$), then a charging station $a$ is added to $C$ at location $\ell_{a}=\ell_{i}^{s}$ ($\ell_{a}=\ell_{i}^{e}$) and the corresponding values of $t_{ij}$, $d_{ij}$ and $p_{ij}$ for $i=a$ or $j=a$ need to be adapted accordingly. Only then the graph is constructed.
@@
Furthermore, fuel cell-electric vehicles can be considered with our methodology if the charging stations $a\in C$ represent hydrogen fueling stations, and the energy consumption rate $\theta$, the charging rate $r$, the maximum energy level $s^{max}$, the minimum energy level $s^{min}$ and $s^{min}_{dep}$ and the minimum time that needs to be available for charging $t^{min}$ are modified to fit for the fuel cell case.
Thus, our approach is universal in the sense that it can be adapted to various zero emission technology settings. Also, the consideration of diesel vehicles is possible with our model analogously to the fuel cell case.
@@ 5 Branch-and-cut algorithm
In order to employ the previously introduced 3-index formulation (1) and 2-index formulation (2) to solve the MDEVSP, one could use one of the many MILP solvers available in the standard configuration. While this is possible for (1), the large number of infeasible path constraints (2f) or connectivity constraints (3) becomes prohibitive for (2).
Thus, we have developed a branch-and-cut algorithm for (2). Branch-and-cut algorithms incorporate the principles of branch-and-bound and pair it with the cutting-plane idea. They start from solving the linear relaxation of the MILP, while considering only a reasonable subset of the original constraints. Typically, constraints of exponential size are excluded. Then, in the course of the algorithm, a separation method is required, which finds violated constraints and adds them in an iterative fashion, until ultimately no original constraints are violated anymore, even though they might not be included explicitly.
@@ 6 Computational experiments
We are now able to present computational results. In the following, first, the attributes of the generated test instances are described in Section 6.1. Then, our obtained results are discussed for instances with a single depot in Section 6.2, a low number of depots (two, three, and four) in Section 6.3, and a high number of depots (six and eight) in Section 6.4.
@@ 6.1 Computational setup and test instances
Everything is implemented in Julia 1.11.1. CPLEX 22.1 is employed for the branch-and-cut algorithms, where for the separation as described in Section 5 we utilize LazyConstraint for separating integer solutions and UserCut for separating fractional solutions. In the separation of the connectivity constraints (3) described in Section 5.2, we use the Boykov-Kolmogorov algorithm within the function maximum_flow() of Julia in order to determine a minimum cut. All experiments were carried out on a Quad-core X5570 Xeon CPU @2.93GHz with a memory of 48 GB.
@@
For our computational experiments, we use a set of MDEVSP benchmark instances generated in a similar way to the generation of MDVSP class A instances in Carpaneto et al., (1989), used by many other researchers (see Gkiotsalitis et al., (2023); Bianco et al., (1994); Fischetti et al., (2001); Forbes et al., (1994); Ribeiro and Soumis, (1994)). The instances of Carpaneto et al., (1989) are not directly applicable to our setting, as they do not involve electric vehicles, which necessitate the inclusion of charging infrastructure in the instance data. Thus, we have created instances as close as possible to Carpaneto et al., (1989) and adapted them to the electric case, as it was done in Gkiotsalitis et al., (2023).
@@
In particular, we consider instances with $|V^{I}|\in\{10,20,30,40,50,60,70,80\}$ service trips, $|K|=|O|=|D|\in\{1,2,3,4,6,8\}$ depots, and $|C|\in\{1,2,3,4,6\}$ charging stations. For each instance, we first determine a number $\nu$ of potential start and end locations of service trips (so-called relief locations) by choosing $\nu$ as uniformly random integer in $\left[|V^{I}|/3,|V^{I}|/2\right]$. Then we choose the $\nu$ relief locations $\ell^{ST}_{1}$, …, $\ell^{ST}_{\nu}$, the depot locations $\ell_{k}$ for each $k\in K$, as well as the charging station locations $\ell_{a}$ for each $a\in C$ randomly distributed in a 60 km by 60 km square in the Euclidean plane using a uniform distribution, resulting in coordinates (latitude and longitude) for each location. Finally, for each service trip $i\in V^{I}$, we decide if it is a short trip (with a probability of $40\%$, representing urban journeys) or a long trip (with probability $60\%$, representing extra-urban journeys that start and end at the same location). Then for each service trip $i\in V^{I}$ we choose both the start and the end location of the service trip $\ell^{s}_{i}$ and $\ell^{e}_{i}$ uniformly at random as one of the $\nu$ relief locations $\ell^{ST}_{1}$, …, $\ell^{ST}_{\nu}$, where we make sure that $\ell^{s}_{i}=\ell^{e}_{i}$ holds for long service trips $i$. From these locations, we compute the distances $d_{ij}$ as the Euclidean distances between the (end) location of $i$ and the (start) location of $j$ for each $i,j\in V^{I}\cup O\cup D\cup C$. For the travel times $t_{ij}$ (in minutes), we assume an average vehicle speed of 60 km per hour, translating to 1 km per minute, so $t_{ij}=d_{ij}$.
@@
For each short service trip $i\in V^{I}$, we generate the start time $s_{i}$ (in minutes since midnight) as a random integer in the interval [420,480] with a probability of 15%, in the interval [480,1020] with a probability of 70% and in the interval [1020,1080] with a probability of 15%. Moreover, we choose the end time $e_{i}$ uniformly at random as integer in the interval $s_{i}+d_{i}+5,s_{i}+d_{i}+40]$, where $d_{i}$ is the Euclidean distance between $\ell^{s}_{i}$ and $\ell^{e}_{i}$. For each long service trip $i\in V^{I}$ we generate the start time $s_{i}$ as a uniform random integer in the interval [300,1200] and the end time $e_{i}$ as a uniform random integer in the interval $s_{i}+180,s_{i}+300]$. For each service trip $i\in V^{I}$, this yields the duration $u_{i}$, from which we compute the energy usage $q_{i}$ as $q_{i}=\theta u_{i}$.
@@
For each depot index $k\in K$, we generate the number of vehicles available as $b_{k}$ as a uniformly random integer in $3+\frac{1}{3|K|}V^{I},3+\frac{1}{2|K|}|V^{I}|]$. The parameters related to vehicles and charging are set to $\theta=1.3$, $s^{max}=1000$, $s^{min}=10$, and $r=50/6$, as done in Gkiotsalitis et al., (2023). Note, that this yields $t^{max}=(s^{max}-s^{min})/r=118.8$. Furthermore, we use $s^{min}_{dep}=0.7s^{max}$ and $t^{min}=s^{max}/100$.
@@
For each instance, we first generate the graph $G=(V,A)$ as described in Section 3.2. As this can be done quite fast (39.63 seconds for the largest instance) we omit the running time for doing so from now on. This yields graphs with a broad range of sizes, ranging from $|V|=29$ and $|A|=172$ for one of the smallest instances with $|V^{I}|=10$, $|K|=1$ and $|C|=1$ up to $|V|=2,187$ and $|A|=22,624$ for one of the largest instances with $|V^{I}|=80$, $|K|=8$ and $|C|=6$. Note that the majority of the nodes of $G$ is charging nodes (for example, $|V^{C}|=2,091$ for the latter instance with $|V|=2,187$).
@@
The graphs $G=(V,A)$ enable us to utilize the 3-index formulation (1) and the 2-index formulation (2) and its variant. We carefully engineered $w_{1}=100,000$, $w_{2}=4,000$ and $w_{3}=1$ as appropriate weights for the weights of the objective functions of (1) and (2) for all our instances, to make sure that we indeed minimize the desired quantities (number of vehicles, number of charging events, energy usage for deadhead trips) in lexicographic order.
@@ 6.2 Results for a single depot
Table 2: Results for a single depot.
$|V^{I}|$ $|C|$ #opt tt(s) gap(%) #nB&B $z^{*}$ $LB_{R}$
10 1 5 0.03 0.00 0.00 473,330.79 472,278.81
10 2 5 0.05 0.00 0.00 475,884.69 475,190.80
10 3 5 0.05 0.00 0.00 555,755.03 555,511.01
20 1 5 0.04 0.00 2.20 866,918.71 866,913.57
20 2 5 0.31 0.00 631.60 850,045.99 845,011.42
20 3 5 0.29 0.00 310.80 765,020.56 761,797.69
30 1 5 1.33 0.00 2,217.80 1,117,568.26 1,104,410.40
30 2 5 0.42 0.00 413.40 1,158,258.26 1,150,531.03
30 3 5 2.80 0.00 2,182.80 1,157,455.10 1,143,326.29
40 1 5 7.05 0.00 4,183.40 1,471,754.78 1,448,886.60
40 2 5 46.27 0.00 38,245.80 1,489,207.54 1,470,915.90
40 3 4 2,169.68 0.04 578,702.00 1,468,933.86 1,450,591.86
50 1 5 109.11 0.00 45,634.00 1,616,374.96 1,578,331.35
50 2 5 647.97 0.00 356,144.00 1,738,233.62 1,699,686.21
50 3 4 2,227.34 0.14 640,487.00 1,659,648.07 1,629,222.31
60 1 3 5,685.70 0.12 1,032,625.60 2,046,328.13 1,998,597.77
60 2 4 2,164.00 0.04 370,454.00 2,210,560.61 2,167,466.49
60 3 4 2,373.70 0.09 512,938.60 1,969,974.84 1,934,643.82
In Table 2 we display results for a single depot, solved with (VI) valid inequalities. In particular, we present the number (#opt) of instances that were solved to optimality (out of 10), the average CPU times ($t$) in seconds, the average percentage gap at the time limit (3 hours) of all instances, the average number of B&B nodes (#nB&B), the average best found objective function value ($z^{*}$) at termination and the average lower bound at the root node ($LB_{R}$).